OpenStudy (anonymous):

A farmer has to transport 20 people and 32 sheep to a market. He can use either Fiats (x) which take 2 people and 1 sheep, or Rolls Royce (y) which take 2 people and 4 sheeps. He must not use more than 15 cars altogether. a) What is the lowest total numbers of cars he could use? b) If it costs $10 to hire each Fiat and $30 for each Rolls Royce, what is the cheapest solution?

OpenStudy (anonymous):

Please help if possible, this is really urgent! Thanks

OpenStudy (anonymous):

It relates to linear programming.

OpenStudy (tylerd):

ok well X = 2p+s Y= 2p+4s aX+bY= 20p+32s a+b=15

OpenStudy (tylerd):

maybe that could get us started

OpenStudy (tylerd):

i guess we should say \[a+b \le 15\]

OpenStudy (tylerd):

@aum

OpenStudy (anonymous):

Oh forgive me but I forgot to add that the answer for a) is 10. I need to know how to get there.

OpenStudy (anonymous):

I should have said that earlier and b) is 250, (4,7).

OpenStudy (tylerd):

so you already solved it?

OpenStudy (anonymous):

no, it is at the back of the textbook

OpenStudy (anonymous):

i have access to answers but not how to solve them.

OpenStudy (tylerd):

ive got no clue about linear programming but for some reason i feel like this involves graphing.

OpenStudy (anonymous):

I believe it does too, that is what linear programming has to do with in my curriculum.

OpenStudy (anonymous):

I guess i just need to know sufficient inequalities to graph them.

OpenStudy (aum):

Assume 'F' Fiats and 'R' Rolls Royces. Each Fiat can transport 2 people and each Rolls can transport 2 people. 2F + 2R >= 20 people F + R >= 10 ---- (1) Each Fiat can transport 1 sheep and each Rolls can transport 4 sheeps. F + 4R >= 32 ---- (2) Minimize F+R. From (1) it looks like the minimum is 10.

OpenStudy (anonymous):

Oh wow, that helps a lot, thank you, now if you can just help with b)

OpenStudy (tylerd):

+fan dude, now im gonna pass out

OpenStudy (anonymous):

hahaha

OpenStudy (aum):

For b) minimize 10F + 30R

OpenStudy (anonymous):

F+3R

OpenStudy (anonymous):

then what do i do?

OpenStudy (aum):

They ask for the cheapest solution so you have to find the minimum value of 10F+30R.

OpenStudy (anonymous):

and for that, i guess i must draw a graph

OpenStudy (aum):

yes. Treat F as y and R as x.

OpenStudy (anonymous):

oh thanks so much, you guys have been a great help

OpenStudy (aum):

You are welcome. What is the book answer for b)?

OpenStudy (anonymous):

I shall take it from here thanks.

OpenStudy (anonymous):

240, (4,7)

OpenStudy (anonymous):

this has proves to be a great website so far, i just started today.

OpenStudy (anonymous):

Thanks once again, I am a fan...literally!

OpenStudy (aum):

You are very welcome. Glad to be able to assist.

OpenStudy (anonymous):

Bye! I am closing now